A plane flies 450 mi/h. It can travel 980 miles with a wind in the same amount of time as it travels 820 miles against the wind. Find the speed of the wind.
You would use T = D/R to solve, correct?
Distance = 980 mi, 820 mi
Time = x
Rate = 450 mi/h

Answer by mananth(16949)

plane speed = 450 mph
wind speed = x mph
wind speed
spee against current 450-x
speed with current 450+ x
Distance against wind 820 miles
Distance with wind 980 miles
t=d/r time against = time with
820/(450-x)=980 /(450+x)
820(450 +x)=980 (450-x)
369000+820x=441000-980 x
980x-820x=72000
1800x=72000
/1800
x=40mph wind speed 0.92